# AFLNet Five Years Later: On Coverage-Guided Protocol Fuzzing

### v1

@@ 6 Recent Progress in Stateful Fuzzing
113      To address this limitation, a series of works have proposed alternative state representations. SGFuzz [28] uses the sequence of values assigned to state variables (identified as enum type variables) to represent the sequence of protocol states corresponding to a message sequence. Similarly, NSFuzz [26] introduces a variable-based state representation to infer states of network protocols. StateAFL [27] infers protocol states by taking snapshots of long-lived memory areas. Utilizing the recent advances in large language models (LLMs), ChatAFL [24] uses LLMs to infer protocol states based on exchanged messages. These approaches effectively address the challenge of state identification in AFLnet.
@@ 7 Reflections and Path Forward
123      Why has the work of AFLNet generated such practical and academic impact in a short period of less than five years? Looking back and reflecting on it, we feel this is because of the sheer dearth of suitable approaches for testing reactive systems, though there exist many approaches for testing sequential transformational systems. Reactive systems are in continuous interaction with the environment by exchanging messages or events between the system and the environment. Thus the “input” to a reactive system is not a single event but rather a sequence of events. Most protocol implementations are reactive systems - instead, the sequence of messages that can be legitimately exchanged is the so-called protocol! Prior to the greybox approach of AFLNet, reactive system validation would typically need to be carried out via stateful blackbox fuzzing approaches or via algorithmic whitebox verification approaches such as model checking. We already mentioned the deficiencies of using stateful blackbox fuzzing approaches - since they involve manual writing of a state model and data model. Moreover, the effectiveness of the stateful blackbox fuzzing approach depends on how complete the manually written state model and data model are.
@@
  • finally after a bug is found, it is reported in the form of a counter-example trace from where a buggy event sequence still needs to be extracted.

### v3

@@ 6 Recent Progress in Stateful Fuzzing
113      To address this limitation, a series of works have proposed alternative state representations. SGFuzz [28] uses the sequence of values assigned to state variables (identified as enum type variables) to represent the sequence of protocol states corresponding to a message sequence. Similarly, NSFuzz [26] introduces a variable-based state representation to infer states of network protocols. StateAFL [27] infers protocol states by taking snapshots of long-lived memory areas. Utilizing the recent advances in large language models (LLMs), ChatAFL [24] uses LLMs to infer states based on exchanged messages. These approaches effectively address the challenge of state identification in AFLnet.
@@ 7 Reflections and Path Forward
123      Why has our work on AFLNet generated such practical and academic impact in a short period of fewer than five years? We can see two reasons: (i) our open science approach and (ii) providing a practical solution to a long-standing problem of validating reactive systems. As for our open science approach, we strongly believe that sound progress in science requires reproducibility and that effective impact in practice requires open source. AFLNet is an excellent case demonstrating the success of our open science approach. Today, it is expected that the tools and artifacts are published together with the paper. Five years ago, it was not common to make prototypes publicly available as open source [47].
124      AFLNet is a practical solution to the long-standing problem of validating reactive systems. Looking back and reflecting on it, we feel this is because of the sheer dearth of suitable approaches for testing reactive systems, though there exist many approaches for testing sequential transformational systems. Reactive systems are in continuous interaction with the environment by exchanging messages or events between the system and the environment. Thus the “input” to a reactive system is not a single event but rather a sequence of events. Most protocol implementations are reactive systems - instead, the sequence of messages that can be legitimately exchanged is the so-called protocol! Prior to the greybox approach of AFLNet, reactive system validation would typically need to be carried out via stateful blackbox fuzzing approaches or via algorithmic whitebox verification approaches such as model checking. We already mentioned the deficiencies of using stateful blackbox fuzzing approaches - since they involve manual writing of a state model and data model. Moreover, the effectiveness of the stateful blackbox fuzzing approach depends on how complete the manually-written state model and data model are.
@@
  • Finally, after a bug is found, it is reported in the form of a counter-example trace from where a buggy event sequence still needs to be extracted.
@@
Acknowledgments
This research is partially supported by the National Research Foundation, Singapore, and Cyber Security Agency of Singapore under its National Cybersecurity R&D Programme (Fuzz Testing NRF-NCR25-Fuzz-0001). Any opinions, findings and conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore, and Cyber Security Agency of Singapore. This research is also partially funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work is supported by ERC grant (Project AT SCALE, 101179366).